OpenStudy (anonymous):

Which is a solution to the equation 5x – y = –3? A. (2, –7) B. (1, 2) C. (–2, 7) D. (–2, –7)

OpenStudy (texaschic101):

time to sub 5x - y = -3 (2,-7)...x = 2 and y = -7 5(2) - (-7) = -3 10 + 7 = 3 17 = 3 (not this one) (1,2)..x = 1 and y = 2 5(1) - 2 = -3 5 - 2 = -3 3 = -3 (nope...not this one) (-2,7)..x = -2 and y = 7 5(-2) - 7 = -3 -10 - 7 = -3 -17 = -3 (nope) (-2,-7)...x = -2 and y = -7 5(-2) - (-7) = -3 -10 + 7 = -3 -3 = -3 (BINGO...it would have to be the last one I tried..lol)